# Wormhole geometries in Einstein-aether theory

### v2

@@ 5.3 The hyperbolic shape function
where the definitions have been used for the sake of brevity. Remember that in this case $r_{0}>0$ must be satisfied due to the conditions (3.2)-(3.4). The region in which the WEC is respected is depicted in Fig. 8. It is evident that by selecting appropriate values within the bound $c_{3}-c_{4}>1$, the wormhole solutions satisfy the WEC at the wormhole throat and at large distances, except in a narrow range near the throat. The same behavior is observed for the NEC for the wormhole solutions in this case.
Figure 8: The WEC and NEC are respected in the blue regions for the wormhole solutions of the EA theory with hyperbolic shape function by choosing appropriate values for the couplings in the bound $c_{3}-c_{4}>0$. In this figure we set $r_{0}=2$ and $e=5$, however we checked varying $e$ and $r_{0}$ does not change this total behavior.

### v3

@@ 5.3 The hyperbolic shape function
where the definitions have been used for the sake of brevity. Remember that in this case $r_{0}>0$ must be satisfied due to the conditions (3.2)-(3.4). The region in which the WEC is respected is depicted in Fig. 8. It is evident that by selecting appropriate values within the bound $c_{3}-c_{4}>0$, the wormhole solutions satisfy the WEC at the wormhole throat and at large distances, except in a narrow range near the throat. The same behavior is observed for the NEC for the wormhole solutions in this case.
Figure 8: The WEC and NEC are respected in the blue regions for the wormhole solutions of the EA theory with hyperbolic shape function by choosing appropriate values for the couplings in the bound $c_{3}-c_{4}>0$. In this figure we set $r_{0}=2.1$ and $e=4$.